OpenStudy (omgthatgirl):

What is the solution of the equation x^2-2x=22 a. x=1+sqrt21 b. x=1+sqrt23 c. x=-1+sqrt21 d. x=-1+sqrt23

OpenStudy (anonymous):

easiest here is to complete the square do you know how to do it?

OpenStudy (anonymous):

if the answer is "no" that is fine, i can show you

OpenStudy (omgthatgirl):

Ok thanks

OpenStudy (anonymous):

what is half of \(2\)?

OpenStudy (omgthatgirl):

1

OpenStudy (anonymous):

right, and \(1^2=1\) so start with \[(x-1)^2=22+1\]

OpenStudy (omgthatgirl):

ok then x-1=22

OpenStudy (anonymous):

then since \(22+1=23\) you have \[(x-1)^2=23\] take the square root of both sides, then add 1 to both sides and you are done

OpenStudy (omgthatgirl):

oh ok

OpenStudy (omgthatgirl):

so the sqrt is 23

OpenStudy (omgthatgirl):

x=1+sqrt23?

OpenStudy (anonymous):

yes \[x-1=\sqrt{23}\]

OpenStudy (omgthatgirl):

negative?

OpenStudy (anonymous):

so \(x=1+\sqrt{23}\) right

OpenStudy (omgthatgirl):

ok thanks :)

OpenStudy (omgthatgirl):

can you help me with another 1?

OpenStudy (anonymous):

sure

OpenStudy (omgthatgirl):

x^2+6x+41=0 a. x=-3+1 sqrt41 b. x=-3+4i sqrt2 c. x=-3+5sqrt2 d. x=3+4i sqrt 2

OpenStudy (anonymous):

lets to this one the same way start with \[x^2+6x=-41\] then complete the square what is half of \(6\)?

OpenStudy (omgthatgirl):

3

OpenStudy (anonymous):

and \(3^2\)?

OpenStudy (omgthatgirl):

9

OpenStudy (anonymous):

ok then we go right from \[x^2+6x=-41\] so \[(x+3)^2=-41+9\] what is \[-41+9\]?

OpenStudy (omgthatgirl):

-32

OpenStudy (anonymous):

right so \[(x+3)^2=-32\] making \[x+3=\pm\sqrt{-32}\]

OpenStudy (omgthatgirl):

ok..

OpenStudy (anonymous):

subtract \(3\) from both sides

OpenStudy (anonymous):

do you know how to rewrite \(\sqrt{-32}\) using \(i\)?

OpenStudy (omgthatgirl):

no..sorry :I

OpenStudy (anonymous):

ok lets do that

OpenStudy (omgthatgirl):

ok

OpenStudy (anonymous):

\[-32=16\times 2\times -1\]si \[\sqrt{-32}=\sqrt{16\times 2\times -1}=\sqrt{16}\sqrt2\sqrt{-1}=4\sqrt2i\]

OpenStudy (omgthatgirl):

ok so we know its 4sqrt2i?

OpenStudy (omgthatgirl):

or 4isqrt2

OpenStudy (anonymous):

right final answer \[x+3=4\sqrt2 i\\ x=-3+4\sqrt2 i\]

OpenStudy (omgthatgirl):

so itts b

OpenStudy (anonymous):

yeah they put the \(i\) in the middle because they don't know what they are doing

OpenStudy (anonymous):

but B is right

OpenStudy (omgthatgirl):

lol ok thanks

OpenStudy (anonymous):

yw